# Fast Transaction Scheduling in Blockchain Sharding

### v1

@@ 3. Technical Preliminaries
A block is a data structure that consists of a set of transactions. The block header contains additional metadata, including the block hash, previous block hash, block sequence, etc. Blockchain is simply a chain of blocks, where a block points to its predecessor (parent) through its hash, which makes them immutable. A blockchain is essentially implemented as a decentralized peer-to-peer ledger where the blockchain is replicated across multiple interconnected nodes.
3.1. Blockchain Sharding Model
@@
We consider blockchain sharding model which is similar to  (Adhikari and Busch, 2023; Hellings and Sadoghi, 2021; Adhikari et al., 2024a, b), consisting of $n$ nodes which are partitioned into $s$ shards $S_{1},S_{2},\dots,S_{s}$ such that $S_{i}\subseteq\{1,\ldots,n\}$, for $i\neq j$, $S_{i}\cap S_{j}=\emptyset$, $n=\sum_{i}|S_{i}|$, and $n_{i}=|S_{i}|$ denotes the number of nodes in shard $S_{i}$. Let $G_{s}=(V,E,w)$ denote a weighed graphs of shards, where $V=\{S_{1},S_{2},\dots,S_{s}\}$, and the edges $E$ correspond to the connections between the shards, and the weight represents the distance between the shards.
@@
Shards communicate with each other via message passing, and here, we are not focusing on optimizing the message size. Moreover, all honest nodes in a shard agree on each message before transmission (e.g. running the PBFT (Castro et al., 1999) consensus algorithm within the shard). Similar to work in (Hellings and Sadoghi, 2021, 2022; Adhikari et al., 2024a, b), we assume that we are given a cluster-sending protocol for reliable and secure communication between shards, satisfying the following properties for sending data $\mathfrak{D}$ from shard $S_{x}$ to shard $S_{y}$:
  • Shard $S_{x}$ sends $\mathfrak{D}$ to $S_{y}$ if there is an agreement among the honest nodes in $S_{x}$ to send $\mathfrak{D}$;
  • All honest nodes in recipient shard $S_{y}$ will receive the same data $\mathfrak{D}$;
  • All non-faulty nodes in sender shard $S_{x}$ receive confirmation of data $\mathfrak{D}$ receipt.
We assume that the above cluster-sending properties between two shards at unit distance are guaranteed to be satisfied within a single communication round. We assume that the communication is synchronous and time is divided into rounds. Moreover, the distance between two shards is measured as the number of rounds required to send a message from one shard to another over the graph $G_{s}$. Note that some of the messages need to propagate through multiple shards in $G_{s}$.
@@
Similar to work in (Adhikari et al., 2024a, b), for secure inter-shard communication, we use a broadcast-based protocol that can operate in a system with Byzantine failures (Hellings and Sadoghi, 2022). For the communication between shard $S_{1}$ and $S_{2}$, shard $S_{1}$ uses a consensus protocol to reach agreement on a value $v$. Then, a set $M_{1}\subseteq S_{1}$ of $f_{1}+1$ nodes in $S_{1}$ and a set $M_{2}\subseteq S_{2}$ of $f_{2}+1$ nodes in $S_{2}$ are chosen, where $f_{i}$ is the number of faulty nodes in shard $S_{i}$. Finally, each node in $M_{1}$ is instructed to broadcast $v$ to all nodes in $M_{2}$. Due to the size of $M_{1}$ and $M_{2}$, at least one non-faulty node in $S_{1}$ will send a value to a non-faulty node in $S_{2}$, which is sufficient to receipt and confirmation of $v$ in $S_{2}$.
@@
Each shard consists of a local blockchain (which is part of the global blockchain) according to the subtransactions it receives and commits. We use $f_{i}$ to represent the number of Byzantine nodes in shard $S_{i}$. To guarantee consensus on the current state of the local blockchain, we assume that every shard executes the PBFT (Castro et al., 1999) consensus algorithm. In order to achieve Byzantine fault tolerance, each shard $S_{i}$ consists of $n_{i}>3f_{i}$ nodes, where $n_{i}$ is the total number of nodes in $S_{i}$.
@@
Suppose we have a set of shared accounts $\mathcal{O}$ (which we also call objects). Similar to previous work in (Hellings and Sadoghi, 2021; Adhikari and Busch, 2023; Adhikari et al., 2024a), we assume that each shard is responsible for a specific subset of the shared objects (accounts). To be more specific, $\mathcal{O}$ is split into disjoint subsets $\mathcal{O}_{1},\ldots,\mathcal{O}_{s}$, where the set of items under the control of shard $S_{i}$ is represented by $\mathcal{O}_{i}$. Every shard $S_{i}$ keeps track of local subtransactions that use its corresponding $\mathcal{O}_{i}$ to access objects.
@@
We assume that shards are connected in a communication graph, and the diameter of the graph is $D$. Moreover, we assume that the distance between the transaction and its accessing objects are in ranges from $1$ to $d\leq D$.
3.2. Transactions and Subtransactions
@@
Let’s consider a set of transactions ${\mathcal{T}}=\{T_{1},T_{2},\ldots\}$ that are distributed across different shards. Let’s say a transaction $T_{i}$ resides in a node $v_{T_{i}}$ within the system. The home shard of $T_{i}$ is the shard containing $v_{T_{i}}$. In this work, we consider the batch transactions that are already in the shard and need to proceed.
@@
Similar to work in (Hellings and Sadoghi, 2021; Adhikari and Busch, 2023; Adhikari et al., 2024a, b), we define a transaction $T_{i}$ as a group of subtransactions $T_{i,a_{1}},\ldots,T_{i,a_{j}}$. Each subtransaction $T_{i,a_{l}}$ accesses objects only in $\mathcal{O}{a_{l}}$ and is associated with shard $S{a_{l}}$. Therefore, each subtransaction $T_{i,a_{l}}$ has a respective destination shard $S_{a_{l}}$. The home shard of $T_{i}$ sends subtransaction $T_{i,a_{l}}$ to shard $S_{a_{l}}$ for processing, where it is appended to the local blockchain of $S_{a_{l}}$. The subtransactions within a transaction $T_{i}$ are independent, meaning they do not conflict and can be processed concurrently. Like previous work in (Hellings and Sadoghi, 2021; Adhikari et al., 2024a, b), each subtransaction $T_{i,a_{l}}$ consists of two parts: (i) a condition check, where it verifies whether a condition of the objects in $\mathcal{O}{a_{l}}$ is satisfied, and (ii) the main action, where it updates the values of the objects in $\mathcal{O}{a_{l}}$.
Example 0.
@@
The home shard of $T_{1}$ sends the subtransactions to their respective destination shards. If the conditions are met (for example, in $T_{1,b}$ if Balen has 700) and the transactions are valid (for example, in $T_{1,r}$ if Ray indeed has 3000 in the account to be deducted), then the destination shards are prepared to commit the subtransactions in the local blockchains, which implies the entire transaction $T_{1}$ implicitly commits as well. However, if any of the conditions in the subtransactions are not met or the subtransactions are invalid, then all the subtransactions must abort (i.e., they are not added to the local blockchains), resulting in the abortion of $T_{1}$ as well.
Definition 0 (Conflict).
@@
4. Centralized Scheduler
We consider a shard graph $G_{s}$ built from $s$ shards, where a set of transactions ${\mathcal{T}}$ and a set of objects ${\mathcal{O}}$ are distributed across the $G_{s}$. In this centralized scheduler, we consider one shard as the leader shard $\ell$, which has global knowledge of the transactions that need to be processed. The leader shard runs a greedy coloring algorithm to schedule transactions in a conflict-free manner. The parameters ($G_{s},{\mathcal{T}},\ell$) are given as input to the algorithm, which then produces a schedule and processes the set of transactions ${\mathcal{T}}$. The details appear in Algorithm 1.
Consider each transaction $T_{i}\in{\mathcal{T}}$ accesses $k$ objects from a set of objects ${\mathcal{O}}$. Suppose $G_{\mathcal{T}}$ is the transactions interference (conflict) graph where each node represents a transaction and an edge between two nodes correspond to a conflict between them. The conflict occurs when the respective transactions try to access the same account (object) $o_{j}\in{\mathcal{O}}$ in the same shard. In Phase 1 of Algorithm 1, our scheduler schedules these transactions using a greedy vertex coloring of $G_{{\mathcal{T}}}$, which assigns execution times to each transaction based on the coloring of $G_{\mathcal{T}}$. A valid coloring of $G_{\mathcal{T}}$ assigns a unique positive integer to each transaction such that two adjacent transactions (conflicting transactions) receive different colors. The colors correspond to the distinct time steps (also called rounds) where the respective transactions are processed for commit. For each color, our algorithm takes four steps to confirm and commit the transaction in each destination shard, which is also shown in Phase $2$ of Algorithm 1. We continue to describe each phase of the algorithm as follows:
@@ 4. Centralized Scheduler
Input : Shard graph $G_{s}$; set of txns ${\mathcal{T}}$ that need to be processed in $G_{s}$; $\ell$ is the leader shard for txns ${\mathcal{T}}$;
@@
7Home shard of each txn $T_{i}\in{\mathcal{T}}$ shares the txn information with the leader shard $\ell$;
@@
9Leader shard $\ell$ constructs $G_{\mathcal{T}}$, and runs a greedy coloring on $G_{\mathcal{T}}$ which returns $\xi$ colors;
10Leader shard $\ell$ informs the home shards of each txn $T_{i}\in{\mathcal{T}}$ about the respective txn colors;
@@
Algorithm 1 CentralizedScheduler ($G_{s}$, ${\mathcal{T}}$, $\ell$)
@@
Phase 1: In this phase, each home shard of transaction $T_{i}\in{\mathcal{T}}$ sends its transaction information to the leader shard $\ell$. Then, the leader shard constructs the transactions conflict graph $G_{\mathcal{T}}$ and runs the greedy coloring algorithm on $G_{\mathcal{T}}$. The greedy coloring algorithm assigns the colors to each transaction of $G_{\mathcal{T}}$ using at most $\xi$ colors. After that, the leader shard informs the home shard of each transaction $T_{i}\in{\mathcal{T}}$ about the respective transaction color.
@@
Phase 2: In this phase, each transaction is processed based on a computed schedule using the greedy coloring in Phase 1. Each home shard knows its transaction committing time and proceeds to determine whether it can be committed or not. Each color in $\xi$ takes $4$ steps for the confirmation (i.e., $4$ communication interactions between home shard and destination shards). So, for each color $clr\in\xi$ in Step 1, each home shard of transaction $T_{i}\in{\mathcal{T}}$ of color $clr$ splits the transaction $T_{i}$ into subtransactions ($T_{i,j}$) and sends them to the respective destination shards parallelly to check whether that transaction (i.e. its subtransactions) can be committed or not. In Step 2, each destination shard $S_{j}$ checks the conditions and validity of the subtransactions $T_{i,j}$ (such as account balance) and sends a commit vote to the home shard of $T_{i}$ if the subtransaction $T_{i,j}$ is valid; otherwise sends an abort vote to home shard of $T_{i}$ if the subtransaction $T_{i,j}$ is invalid. In Step 3, the home shard of $T_{i}$ collects the votes, and if it receives all the commit votes for a transaction $T_{i}$ from its destination shards, then it sends the confirmed commit message to the respective destination shards. Otherwise, if the home shard of $T_{i}$ receives any abort vote, then it will send a confirmed abort message to all respective destination shards. In Step 4, each destination shard either commits the subtransactions $T_{i,j}$ and appends them to a local ledger or aborts the subtransactions $T_{i,j}$ according to the message received from the home shard of $T_{i}$.
4.1. Analysis for General Graph
Suppose we run Algorithm 1 for the general graph whose diameter is $D$ and the distance between the home shard (transaction) and it’s accessing objects (destination shards) ranges from $1$ to $d$, where $d\leq D$. In the following section, we provide an analysis of such a case.
Lemma 0 (Lower Bound).

### v3

@@ III Technical Preliminaries
A block is a data structure that consists of a set of transactions. The block header contains additional metadata, including the block hash, previous block hash, block sequence, etc. Blockchain is simply a chain of blocks, where a block points to its predecessor (parent) through its hash, which makes blocks immutable. A blockchain is essentially implemented as a decentralized peer-to-peer ledger where the blockchain is replicated across multiple interconnected nodes.
Blockchain Sharding Model
@@
We consider blockchain sharding model which is similar to  [15, 22, 16], consisting of $n$ nodes which are partitioned into $s$ shards $S_{1},S_{2},\dots,S_{s}$ such that $S_{i}\subseteq\{1,\ldots,n\}$, for $i\neq j$, $S_{i}\cap S_{j}=\emptyset$, $n=\sum_{i}|S_{i}|$, and $n_{i}=|S_{i}|$ denotes the number of nodes in shard $S_{i}$.
Let $G_{s}=(V,E,w)$ denote a weighted graph of shards, where $V=\{S_{1},S_{2},\dots,S_{s}\}$, and the edges $E$ correspond to the connections between the shards, and the weight represents the distance between the shards. We assume that $G_{s}$ is a clique, namely, any two shards can communicate with each other. The communication delay between two shards is expressed by the edge weight connecting them.
@@
Shards communicate with each other via message passing, and here, we are not focusing on optimizing the message size. Moreover, all honest nodes in a shard agree on each message before transmission (e.g. running the PBFT [25] consensus algorithm within the shard). Similar to [22, 47, 16], we use a cluster-sending protocol for reliable and secure communication between shards, satisfying the following properties: (1) Shard $S_{x}$ sends data $\mathfrak{D}$ to shard $S_{y}$ if there is an agreement among the honest nodes in $S_{x}$ to send $\mathfrak{D}$. (2) All honest nodes in recipient shard $S_{y}$ will receive the same data $\mathfrak{D}$. (3) All non-faulty nodes in sender shard $S_{x}$ receive confirmation of data $\mathfrak{D}$ receipt.
@@
Similar to [16], for secure inter-shard communication, we use a broadcast-based protocol that can operate in a system with Byzantine failures [47]. For the communication between shards $S_{1}$ and $S_{2}$, shard $S_{1}$ uses a consensus protocol to reach agreement on a value $v$. Then, a set $M_{1}\subseteq S_{1}$ of $f_{1}+1$ nodes in $S_{1}$ and a set $M_{2}\subseteq S_{2}$ of $f_{2}+1$ nodes in $S_{2}$ are chosen, where $f_{i}$ is the number of faulty nodes in shard $S_{i}$. Finally, each node in $M_{1}$ is instructed to broadcast $v$ to all nodes in $M_{2}$. Due to the choices of $M_{1}$ and $M_{2}$, it is guaranteed that at least one non-faulty node in $S_{1}$ will send a value to a non-faulty node in $S_{2}$, which is sufficient to receipt and confirmation of $v$ in $S_{2}$.
We assume communication is synchronous and time is divided into rounds. A round is the time it takes to reach a consensus within a shard. A round is also the time to send a message from one shard to another at a unit distance. The above cluster-sending properties between two shards at unit distance are guaranteed to be satisfied within a single communication round. Moreover, the distance between two shards is measured as the number of rounds required to send a message from one shard to another over the graph $G_{s}$.
@@
Each shard consists of a local blockchain (which is part of the global blockchain) according to the subtransactions it receives and commits. We use $f_{i}$ to represent the number of Byzantine nodes in shard $S_{i}$. To guarantee consensus on the current state of the local blockchain, we assume that every shard executes the PBFT [25] consensus algorithm. In order to achieve Byzantine fault tolerance, similar to previous works [22, 16, 15], we assume each shard $S_{i}$ consists of $n_{i}>3f_{i}$ nodes, where $n_{i}$ is the total number of nodes in $S_{i}$.
@@
Suppose we have a set of shared accounts $\mathcal{O}$ (which we also call shared objects). Similar to previous works in [22, 15, 16], we assume that each shard is responsible for a subset of the shared objects (accounts). Thus, $\mathcal{O}$ is partitioned into disjoint subsets $\mathcal{O}_{1},\ldots,\mathcal{O}_{s}$, where $\mathcal{O}_{i}$ is the set of objects handled by shard $S_{i}$.
@@
Let $D$ denote the diameter of graph $G_{s}$. We assume that the distance between the transaction and its accessing objects ranges between $1$ to $d$, where $d\leq D$.
Transactions and Subtransactions
@@
Let’s consider a set of transactions ${\mathcal{T}}=\{T_{1},T_{2},\ldots\}$ that are distributed across different shards. Let $S(T_{i})$ denote the home shard of transaction $T_{i}$, which is the shard responsible for processing $T_{i}$. The home shard may receive $T_{i}$ from one of the IoT or mobile devices associated with the shard. In this work, we consider batch transactions that are appear in the home shards and need to be processed.
@@
As in [22, 15, 16], a transaction $T_{i}$ consists of subtransactions $T_{i,a_{1}},\ldots,T_{i,a_{j}}$, such that each subtransaction $T_{i,a_{l}}$ accesses objects only in $\mathcal{O}{a_{l}}$ associated with shard $S{a_{l}}$. Thus, each subtransaction $T_{i,a_{l}}$ has a respective destination shard $S_{a_{l}}$. The home shard of $T_{i}$ sends subtransaction $T_{i,a_{l}}$ to shard $S_{a_{l}}$ for processing, where it is appended to the local blockchain of $S_{a_{l}}$. The subtransactions within a transaction $T_{i}$ are independent, meaning they do not conflict and can be processed concurrently. Like previous work in [22, 16], each subtransaction $T_{i,a_{l}}$ consists of two parts: (i) a condition check, where it verifies whether a condition of the objects in $\mathcal{O}{a_{l}}$ is satisfied, and (ii) the main action, where it updates the values of the objects in $\mathcal{O}{a_{l}}$.
Example 1
@@
The home shard of $T_{1}$ sends the subtransactions to their respective destination shards. If the conditions are met and the transactions are valid, then the destination shards commit the subtransactions in the local blockchains, which implies the entire transaction $T_{1}$ implicitly commits as well. However, if any of the conditions are not met, they are not added to the local blockchains, resulting in the abortion of $T_{1}$ as well.
Definition 1 (Conflict)
@@
IV Centralized Scheduler
We consider a shard graph $G_{s}$ built from $s$ shards, where a set of transactions ${\mathcal{T}}$ and a set of objects ${\mathcal{O}}$ are distributed across the $G_{s}$. In this centralized scheduler, we consider one shard as the leader shard $S_{\ell}$, which is responsible to create a schedule for all the transactions. The home shards of the various transactions send the transaction information to the leader shard $S_{\ell}$ which it then runs a greedy coloring algorithm to schedule the transactions in a conflict-free manner. The details appear in Algorithm 1.
Consider each transaction $T_{i}\in{\mathcal{T}}$ accesses at most $k$ objects from a set of objects ${\mathcal{O}}$. Suppose $G_{\mathcal{T}}$ is the transactions interference (conflict) graph where each node represents a transaction and an edge between two nodes corresponds to a conflict between them. Recall that a conflict occurs when the respective transactions try to access the same account (object) $o_{j}\in{\mathcal{O}}$ in the same shard.
Algorithm 1 uses two phases. In Phase 1, it performs a greedy vertex coloring of $G_{{\mathcal{T}}}$, which assigns execution times to each transaction based on their respective colors. A valid coloring of $G_{\mathcal{T}}$ assigns a unique positive integer to each transaction such that two adjacent transactions (conflicting transactions) receive different colors. The colors correspond to the distinct time steps (also called rounds) where the transactions are processed for commit. In Phase 2, our algorithm takes four steps for each assigned color to confirm and commit the transactions with that color in each destination shard. We continue to describe each phase of the algorithm.
@@ IV Centralized Scheduler
Input : Shard graph $G_{s}$; set of txns ${\mathcal{T}}$ that need to be processed in $G_{s}$; $S_{\ell}$ is the leader shard for txns ${\mathcal{T}}$;
@@
7Home shard of each txn $T_{i}\in{\mathcal{T}}$ shares the txn information with the leader shard $S_{\ell}$;
@@
9Leader shard $S_{\ell}$ constructs $G_{\mathcal{T}}$, and runs a greedy coloring on $G_{\mathcal{T}}$ which returns $\xi$ colors;
10Leader shard $S_{\ell}$ informs the home shards of each txn $T_{i}\in{\mathcal{T}}$ about the respective txn colors;
@@
Algorithm 1 CentralizedScheduler ($G_{s}$, ${\mathcal{T}}$, $S_{\ell}$)
@@
Phase 1: In this phase, the home shard of each transaction $T_{i}\in{\mathcal{T}}$ sends its transaction information to the leader shard $S_{\ell}$. Then, the leader shard constructs the transactions conflict graph $G_{\mathcal{T}}$ and runs the greedy coloring algorithm on $G_{\mathcal{T}}$. The greedy coloring algorithm assigns the colors to each transaction of $G_{\mathcal{T}}$ using at most $\xi$ colors. After that, the leader shard informs the home shard of each transaction $T_{i}\in{\mathcal{T}}$ about the respective transaction color.
@@
Phase 2: In this phase, each transaction is processed based on a computed schedule using the greedy coloring in Phase 1. Each home shard knows its transaction committing time and proceeds to determine whether it can be committed or not. Each of the $\xi$ colors takes $4$ steps for the confirmation (i.e., $4$ communication interactions between home shard and destination shards). So, for each color $clr\in\xi$ in Step 1, each home shard of transaction $T_{i}\in{\mathcal{T}}$ of color $clr$ splits the transaction $T_{i}$ into subtransactions ($T_{i,j}$) and sends them to the respective destination shards parallelly to check whether that transaction (i.e. its subtransactions) can be committed or not. In Step 2, each destination shard $S_{j}$ checks the conditions and validity of the subtransactions $T_{i,j}$ (such as account balance) and sends a commit vote to the home shard of $T_{i}$ if the subtransaction $T_{i,j}$ is valid; otherwise sends an abort vote to home shard of $T_{i}$ if the subtransaction $T_{i,j}$ is invalid. In Step 3, the home shard of $T_{i}$ collects the votes, and if it receives all the commit votes for a transaction $T_{i}$ from its destination shards, then it sends the confirmed commit message to the respective destination shards. Otherwise, if the home shard of $T_{i}$ receives any abort vote, then it will send a confirmed abort message to all respective destination shards. In Step 4, each destination shard either commits the subtransactions $T_{i,j}$ and appends them to a local ledger or aborts the subtransactions according to the message received from the home shard of $T_{i}$.
Figure 2: Simple representation of transaction processing using Algorithm 1, where (a) represents four shards, each having one pending transaction to process. Figure (b) and (c) represent Phase 1, and Figure (d) represents Phase 2 of the Algorithm 1.
Example of Transaction Processing Using Algorithm 1
Consider four shards, $S_{1}$ to $S_{4}$, each with one pending transaction: $T_{1}$ to $T_{4}$ (shown in Figure 2 (a)). The transactions access accounts on different shards as follows: $T_{1}$ accesses $S_{2}$ and $S_{3}$, $T_{2}$ accesses $S_{2}$ and $S_{4}$, $T_{3}$ accesses $S_{1}$ and $S_{3}$, and $T_{4}$ accesses $S_{3}$ and $S_{4}$. Since each shard holds one account, any transactions that access the same shard will conflict and must be scheduled in different rounds.
Consider shard $S_{1}$ acts as the leader ($S_{\ell}$). During Phase 1, each shard sends its transaction information to the $S_{\ell}$, which constructs a transaction conflict graph $G_{\mathcal{T}}$ (as shown in Figure  2 (b)). Using a greedy coloring algorithm, the leader assigns colors to the transactions based on conflicts: $T_{1}$ receives blue, $T_{2}$ and $T_{3}$ both receive green, and $T_{4}$ receives orange (shown in Figure 2 (c)). The first two rounds are used to share transaction information and to color. Then, the blue transaction $T_{1}$ executes in rounds $3-6$ (i.e., start at round $3$ and completed by $6$ as each color takes $4$ rounds), the green transactions ($T_{2}$, $T_{3}$) in rounds $7-10$, and the orange transaction ($T_{4}$) in rounds $11-14$. In Phase 2, each home shard splits its transaction into subtransactions and sends them to the destination shards for processing based on the assigned rounds and colors. For instance, $S_{1}$ splits $T_{1}$ into $T_{1,2}$ and $T_{1,3}$, sending them to $S_{2}$ and $S_{3}$, respectively (shown in Figure 2 (d)). Transactions are executed in their designated rounds, with all transactions completed by round 14.
In the following, we analyze the time required to process transactions in Algorithm 1, focusing on the period after the leader shard has determined the schedule for all transactions. More specifically, we omit the cost of Phase 1 in this analysis. If the shard is in a clique graph with uniform distances between shards, this cost remains constant. However, if the distance between a home shard and the leader shard is $d$, an additional cost proportional to $d$ is incurred.
IV-A Analysis for General Graph
Suppose we run Algorithm 1 for general graph $G_{s}$. We have the following results.
Lemma 1 (Lower Bound)